# Predictive Planner for Autonomous Driving with Consistency Models

### v2

@@ VI-B3 Ablation Study on ADMM-inspired Guided Sampling
We compare our ADMM-inspired method (ADMM for short) in Eq. (7) with the vanilla gradient descent (GD for short) in Eq. (6), which simultaneously optimizes all planning constraints. To illustrate the advantage of our method in hyperparameter tuning, we sweep through all combinations of step sizes ($\alpha=1e^{-4},1e^{-5},1e^{-6}$) across three constraints (27 combinations total), using 50, 100, and 150 gradient steps over 200 scenarios. TABLE III shows that GD results in significantly higher violations than ADMM on average; e.g., at 100 steps, GD has 9% higher goal-reaching, 35% higher acceleration, and 8% higher angular speed violations. The lowest sums of violations achieved are 1.45, 1.16, and 0.71 for 50, 100, and 150 steps, respectively. Increasing gradient steps reduces constraint violations but with computational tradeoffs.

### v3

@@ VI-B3 Ablation Study on ADMM-inspired Guided Sampling
We compare our ADMM-inspired method (ADMM for short) in Eq. (7) with vanilla gradient descent (GD for short) in Eq. (6), which simultaneously optimizes all planning constraints. To show the advantage of our method in hyperparameter tuning, we sweep through all combinations of step sizes ($\alpha=1e^{-4},1e^{-5},1e^{-6}$) across three constraints (27 combinations total), using 50, 100, and 150 gradient steps over 200 scenarios. TABLE III shows that GD results in significantly higher violations than ADMM on average; e.g., at 100 steps, GD has 9% higher goal-reaching, 35% higher acceleration, and 8% higher angular speed violations. With increasing gradient steps of 50, 100, and 150 steps, we obtain lower sums of violation as 1.45, 1.16, and 0.71, but with more computational tradeoffs.
@@
VI-B4 Computational Time
With the same encoder, the computational time for diffusion-based methods directly depends on the number of sampling steps, shown in Table I. Diffusion-4 and DDIM have the same sampling step as our model, but fail to match our performance in trajectory generation across all metrics in Table I and Table II. The guided sampling, introducing a little computational overhead ($\approx 15\%$), provides higher quality trajectory generation and enables controllable planning. Even Diffusion-10, despite its more sampling steps, cannot achieve a comparable performance of trajectory generation to our approach in Table II.